OpenStudy (anonymous):

Let u = <-5, -3>, v = <-6, -1>. Find -3u + 5v. @hero @dan815 @nincompoop @mathstudent55 @wio @radar @undeadknight26

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

multiply both entries in \(u\) by \(-3\) multiply both entries in \(v\) by \(5\) then add

OpenStudy (anonymous):

huh? @misty1212

OpenStudy (misty1212):

i don't really know another way to say it

OpenStudy (misty1212):

u has two numbers right?

OpenStudy (anonymous):

can u show me? @misty1212

OpenStudy (misty1212):

multiply each of them by \(-3\)

OpenStudy (misty1212):

\[u=<-5,-3>\\ -3u=-3<-5,-3>=<15,9>\]

OpenStudy (misty1212):

in other words, multiply each entry by \(-3\)

OpenStudy (misty1212):

repeat to find \[5v=5<-6,-1>\]

OpenStudy (anonymous):

so its <-30,-5> @misty1212

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

then add

OpenStudy (anonymous):

can u help me with another? @misty1212

OpenStudy (misty1212):

did you add them yet?

OpenStudy (anonymous):

yes i got <-15,4> @misty1212

OpenStudy (misty1212):

oops yes you are right sorry

OpenStudy (anonymous):

can u help me with another? @misty1212

OpenStudy (misty1212):

ok sure

OpenStudy (anonymous):

Find a • b. a = 4i + 3j, b = -4i + 4j -4 28 <-16, 12> <0, 7>

OpenStudy (misty1212):

is that a dot between them?

OpenStudy (anonymous):

i think it times

OpenStudy (misty1212):

i think it means the "dot product"

OpenStudy (misty1212):

\[<4,3>\cdot<-4,4>=4\times (-4)+3\times 4\]

OpenStudy (misty1212):

it is a number, not a vector

OpenStudy (misty1212):

\[-16+12=-4\] in other words

OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -7, an = 2 • an-1 -7, -14, -12, -10, -8, -6 0, 2, -14, -12, -10, -8 -7, -14, -28, -56, -112, -224 -14, -28, -56, -112, -224, -448

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

start with \(-7\)

OpenStudy (misty1212):

then multiply it by 2 to get the next number

OpenStudy (misty1212):

\[-7,-7\times 2,-7\times 2\times 2...\]

OpenStudy (misty1212):

so this one \[ -7, -14, -28, -56, -112, -224\]

OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 4) and a directrix at y = -4. y = 1 divided by 16x2 y2 = 16x y2 = 4x y = 1 divided by 4x2 @misty1212

OpenStudy (misty1212):

i would go with the first one want to check?

OpenStudy (misty1212):

yup looks good

OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 3, -12, 48, -192, ... an = 3 • 4n - 1 an = 3 • 4n + 1 an = 3 • (-4)n - 1 an = 3 • (-4)n @misty1212

OpenStudy (misty1212):

see how the signs change?

OpenStudy (anonymous):

yes by -4 @misty1212

OpenStudy (misty1212):

so only \[a_n=3(-4)^{n-1}\] makes any sense

OpenStudy (anonymous):

one last one? @misty1212

OpenStudy (misty1212):

kk

OpenStudy (anonymous):

Use graphs and tables to find the limit and identify any vertical asymptotes of the function.

OpenStudy (misty1212):

? vertical asymptote is where the denominator is zero, so \(x=5\)

OpenStudy (anonymous):

what is the limit @misty1212

OpenStudy (misty1212):

ooh i see

OpenStudy (misty1212):

limit is infinity \(\infty\)

OpenStudy (misty1212):

oh no oops

OpenStudy (misty1212):

try \(-\infty\)